OpenStudy (anonymous):

How do I find slope of 2 coordinates? Helpl me plz! I have no direct question. I just want to see an example please.

OpenStudy (aravindg):

\[slope=\dfrac{y_2-y_1}{x_2-x_1}\]

OpenStudy (whpalmer4):

If you have two points, (x1, y1) and (x2, y2) (order doesn't matter) \[ m = \frac{y_2-y_1}{x_2-x_1}\] where m is the slope. Proof that order doesn't matter: \[ m = \frac{y_2-y_1}{x_2-x_1} = \frac{-1}{-1} * \frac{(y_2-y_1)}{(x_2-x_1)} = \frac{y_1 - y_2}{x_1 - x_2}\] You do have to be consistent that you use the same order in the numerator and denominator, however!

OpenStudy (anonymous):

So say if I input (3,4) and (6,2) 4-2 --- = 2/-3 3-6 So is 2/-3 the slope? btw how did you do the fancy writing.

OpenStudy (anonymous):

Is my slope correct guys?

OpenStudy (aravindg):

yes

OpenStudy (aravindg):

right :)

OpenStudy (whpalmer4):

Yes, that's the correct slope. It's negative, because as you go to the right on the x-axis, the y-value decreases. It's < 1 in magnitude, because if you move 1 unit on the x-axis, the change in the y-axis is less than 1 unit.

OpenStudy (anonymous):

okay thx guys!

OpenStudy (whpalmer4):

slope = 1 -> line goes up to the right at a 45 degree angle slope = -1 -> line goes up to the left at a 45 degree angle, and the two lines drawn together (slope 1 and -1) are perpendicular and make a nice symmetrical X

OpenStudy (whpalmer4):

(indeed, the way to find a line perpendicular to another is to find the slope of the known line, then find another slope such that the product of the two is -1)

OpenStudy (whpalmer4):

Your line has a slope of -2/3, so the perpendicular line would be -1/(-2/3) = 3/2.